12 : m= 75 : 40
I think you multiply 12 by a number and then divide it by another, but I'm not sure what number to do first.

Answer by rothauserc(4718)

12 : m= 75 : 40 = 12/m = 75/40
cross multiply the fractions
12*40 = 75m
75m = 480
m = 6.4